OpenStudy (anonymous):

solve for x. 5/2x+3 = 4/x x = -1 x = -4 x = 4

OpenStudy (anonymous):

is it \[\frac{5}{2x+3}=\frac{4}{x}\]?

OpenStudy (anonymous):

if so then "cross multiply" and solve \[5x=4(2x+3)\] you good from there?

OpenStudy (anonymous):

x = -4 @Gabriella77

OpenStudy (anonymous):

I could have gotten it but thanks